OpenStudy (anonymous):

derivative of sin^2(x)/x

OpenStudy (yttrium):

Apply quotient rule. Do you know it?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

Actual question is F(t)=sin^2(x)/x then what is F'(t)?

OpenStudy (yttrium):

\[\frac{ vdu-udv }{ v^2 }\] That what you are going to apply. Do you understand this pattern?

OpenStudy (anonymous):

F(t)=sin^2(x)/x = sin^2(x) * x^(-1) Now, use product rule ( product rule and quotient rule are somewhat same) Product rule is : d(f(x) * g(x))/dx = f(x) * g'(x) + f'(x) * g(x)

OpenStudy (anonymous):

hmm i would not recommend that

OpenStudy (yttrium):

what @satellite73

OpenStudy (anonymous):

i would not recommend using the product rule don't be scared of the quotient rule

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\sin^2(x), f'(x)=2\sin(x)\cos(x), g(x)=x, g'(x)=1\]

OpenStudy (anonymous):

Ya i do understand this formula @Yttrium But then to i'm new for it. So how to solve next u=sin^2x and v=t t(d/dt sin^2 t)-sin^2t(1)/t^2 right?? @Yttrium derivative of sin^2 (x) will be what?

OpenStudy (anonymous):

you need the chain rule for that

OpenStudy (anonymous):

Now what is chain rule?? :o

OpenStudy (anonymous):

you have a composite function the square of the sine of \(x\)

OpenStudy (anonymous):

\[\left(f\circ g\right)'=f'(g)g'\]

OpenStudy (anonymous):

in this case \(f(x)=x^2\) and \(g(x)=\sin(x)\) so \(f\circ g(x)=f(\sin(x))=[\sin(x)]^2=\sin^2(x)\)

OpenStudy (anonymous):

so what would be the final answer?

OpenStudy (anonymous):

i wrote it above

OpenStudy (anonymous):

the derivative of something squared is two times something, times the derivative of something the derivative of sine squared is two times sine, times the derivative of sine the derivative of sine squared, is two times sine, times cosine

OpenStudy (anonymous):

Ohh that is so confusing but i understood. :) @Satellite73 So just check F(t)= sin^2(x)/x F'(t)={x[2(sin x)(cos x)]-sin^2 x}/x^2

OpenStudy (yttrium):

@Swara , you got it right! Just continue it to arrive at an answer. :))